# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v2

@@ 4.1.2 Application to MAAPS Data
The ExpertTest* is framed in terms of individual actions at a single timestep $t$ affecting outcomes at $t+1$, so to the multistage, multi-intervention nature of collegiate advising complicates variable definition. We run multiple different variants of the ExpertTest* to accommodate these complexities. Below we describe how we define $\mathbf{X}_{t}$, $A_{t}$, $Y_{t+1}$, and $n$ for these tests. All test parameters and variables, as well as details of how we chose test parameters $L$ and $K$, are summarized in Appendix A.3.3.
@@
For all the tests, we take $\textbf{X}_{t}$, the algorithmic information, as students’ attributes at baseline excluding their race and ethnicity, since Georgia State University excludes this information from all its early warning systems [60]. Following our data cleaning process (§2.1.3), the population consists of 424 students having 3,507 meetings over four years (Figure 5).
The test requires matching students with approximately identical algorithmically-available features. We pair students greedily using a Euclidean distance function applied to fixed-range scaled features (Figure 6).We select the number of pairs for each test, $L$, to yield an acceptable tradeoff between test power, which increases in $L$, and test error, which increases in match distance (Appendix A.3.3).
@@
As discussed in §3.3.3, for actions to assist at-risk students where the risk factors are not encoded in $\mathbf{X}_{t}$, an action that supports students might be negatively associated with outcomes. Thus, we also run each test on $A_{t}^{i-}=|1-A_{t}^{i+}|$, which detects cases in which not taking the action predicts positive outcomes, or equivalently, taking the action predicts negative outcomes.
@@
Of the three outcomes we examine, persistence, graduation and GPA (at various timescales), all but GPA are naturally binary; we binarize the GPA outcome by testing for GPA above the mean of 3.5, testing more thresholds for robustness. [^22] We note that the MSE here differs from its normal interpretation—whereas MSE is normally a comparison of predictions of an outcome to that outcome, here we look at whether actions predict binary outcomes. Future work might explore alternate test statistics or conditional independence tests from Alur et al. [4].
@@
    Semesterly tests: Among enrolled students in each semester (Figure 5), we look at the result of the application of each intervention $i$ in a meeting that semester on term GPA or persistence to the next term (for spring terms, persistence is defined as enrollment/graduated status in either the summer or fall). In these tests, students’ semester of enrollment is added as a feature to their baseline attributes $\mathbf{X}_{t}$, and cross-semester matches of students are prohibited.
@@
4.2 Results and Discussion
4.2.1 Results
For the ExpertTest* on semesterly timescales, four interventions demonstrate significant conditional dependence with GPA $\geq 3.5$ (Table 1) and one intervention demonstrates significant dependence with persistence (Table 2). Significance levels for all 20 interventions are given in Appendix A.3.5, alongside robustness checks demonstrating the consistency of the results to inclusion of potential transfer students, choice of GPA threshold, and exclusion of students assigned advisor 1.
These five significant interventions comprise 5,615 of the 8,385 total interventions taken by advisors following our data cleaning process. Accordingly, we estimate that 67%, or two-thirds, of interventions taken by MAAPS advisors are potentially expertly targeted. We provide this numerical estimate of the pervasiveness of significant interventions, with the caveat that a significant ExpertTest* result is necessary but not sufficient for expertly targeted action to be present. We address this limitation in §4.2.3 and §5, which provides additional qualitative evidence of expert targeting.
Two of these significance results have a positive direction, indicated by a (+)(+)( + ) in Tables 1 and 2. These correspond to tests on $A_{t}^{i+}$; the $p$-value for positive direction results should be interpreted as the probability that actions taken by advisor that cannot be explained by administrative data predict a positive outcome for the student (high GPA or next-term enrollment) just by chance. Three of the significant interventions have a negative direction, corresponding to tests on $A_{t}^{i-}$ and notated by ()(-)( - ), with $p$-values in these cases interpreted as the probability that actions taken by advisors that are unexplained by administrative data predict negative outcomes (low term GPA or failure to enroll next term) just by chance.
We note that the tests of first-year interventions on a four-year timescale do not meet the significance threshold (see results in Appendix A.3.5). We suspect this is due to the long time range of the hypothesis: Advisors’ actions in the first year are unlikely to contain detectable signal about outcomes three years later.
Semesterly tests on high GPA: top five interventions
@@ 4.2.1 Results
Table 1: Top five most significant interventions for the ExpertTest* with $A^{i}_{t}$ indicating application of intervention $i$ during a given term, and $Y_{t+1}$ as that term’s GPA $\geq 3.5$ (the mean student GPA).
Semesterly tests on persistence: top five interventions
@@
Table 2: Top five most significant interventions for the ExpertTest* with $A^{i}_{t}$ indicating application of intervention $i$ during a given term, and $Y_{t+1}$ as persistence to the next term.
4.2.2 Discussion: Effect Sizes and Significance
Interestingly, there is a higher number of significant interventions for high term GPA than for persistence. We hypothesize three reasons for this trend. First, the power of the ExpertTest* is decreased for outcomes with high or low base rates, and persistence to the next semester is relatively common (Appendix A.3.4). Second, failure to persist is a less straightforward negative signal than a low GPA. For example, the comment data indicates a student might not re-enroll because they transferred to a school that better fits their needs, took a break to pursue a valuable opportunity outside of school, or faced family circumstances beyond their control. As a result, high-quality advisor support may not have a unilaterally positive effect on persistence. Third, the effects of actions advisors take to promote persistence might only surface over longer time ranges, whereas semester GPA can easily be impacted by advisors’ recommendations in a particular term (like recommendations to take particular courses).
@@ 4.2.2 Discussion: Effect Sizes and Significance
Additionally, for all outcomes, it is possible that additional interventions meet the necessary condition for expertise but remain undetected. Just like outcomes, the ExpertTest* may not identify dependence for interventions with more extremal base rates—for instance, very rare interventions. This is because, for rare interventions, there are fewer “swaps” of algorithmically-indistinguishable individuals’ intervention values that can affect the mean squared error of a synthetic dataset, as most individuals’ intervention values are already equal to 0. Accordingly, Appendix A.3.4 shows that rarer interventions in the MAAPS dataset have fewer swaps between algorithmically-indistinguishable pairs that could change the MSE.
4.2.3 Discussion: Evidence of Expertise
@@
We recall that the conditional dependence of actions and outcomes is only a necessary condition for expertise and that the test results do not confirm the five significant interventions are truly expertly targeted according to Definition 3.1. Expertly targeted action is in general not identifiable from observational data, as a consequence of Proposition 3.1. In other words, the conditional dependence of outcomes and actions can have alternative explanations than the use of non-algorithmic information, as our theoretical framework (§3.3) clarifies.
The set of interventions that are statistically significant raises three key questions to guide our subsequent qualitative analysis. Each of these questions corresponds to a particular concern about the mechanism that leads to conditional dependence, as discussed in §3.3.4: (A) non-algorithmic targeting, (B) supporting at-risk students and (C) nuisance confounding.
(A) Non-Algorithmic Targeting.
In the case of a significant result (actions predictive of outcomes in either direction), is there evidence that the interventions are non-algorithmically targeted? For instance, do advisors use non-algorithmic information about students to customize the content of students’ schedules (Intervention 17)? This seems plausible, since many factors informing schedule planning, like particular student interests and timing contingencies, could be difficult to predict from basic administrative data about them.
(B) Supporting At-Risk Students.
In the case the intervention is predictive of negative outcomes, is there evidence that advisors are targeting to at-risk students? Most of the significant interventions predict negative, rather than positive, outcomes. If advisors are not targeting on risk factors (and there is no other hidden confounding), then it is possible these actions of advisors hurt student outcomes. Since the interventions that negatively predict outcomes signal that the student might be in trouble (like being advised about degree progression issues or advised to come in for an appointment), we hypothesize that advisors might be targeting based on algorithmically-invisible risk factors.
(C) Nuisance Confounding.
Could any significant results be driven by nuisance confounding? The association of being advised about recommended schedule (Intervention 17) with persistence suggests some interventions might be confounded by nuisance factors. It may be that students who are likely to persist to next semester are also likely to ask their advisors for schedule advice in the first place, introducing dependence between schedule recommendations and outcomes that is not due to targeting.
We turn to qualitative analysis of the advisors’ logs of students interactions to answer these three questions.
@@ 5 Qualitative Analysis: Evidence for Non-Algorithmic Targeted Action
As described in §3 and §4, the quantitative observational data allows us to test for a necessary condition (conditional dependence) for expertly targeted actions, but does not allow us to distinguish expertly targeted actions from other mechanisms that may also be driving the conditional dependence. However, we have access to a rich dataset of qualitative data on advisors’ notes with their students that sheds light on these mechanisms.
We conduct a two-stage qualitative analysis to better understand advisor discretion and intervention practices, complementing our quantitative findings. The first stage involves open coding to identify themes across all advisor comments, to familiarize ourselves with the data and its limitations. The second stage focuses on a targeted analysis of comments linked to statistically significant interventions from the ExpertTest* to investigate potential evidence of expert targeting. In this section, we describe the methods for our qualitative analysis in §5.1 and how they complement our quantitative findings about advisor expertise in §5.2.
@@ 5.1.1 Exploratory Coding
In the first stage of our qualitative analysis, we conduct exploratory coding on all advisor meetings comments to understand the nature of advisor-student interactions as documented by those comments. Following removal of students due to our data cleaning (§2.1.3), we code a total of 1,594 comments left by advisors on meetings with students in the treatment arm. Using the Atlas.ti software, we apply both deductive codes based on our research questions and inductive codes that emerged from the data. To determine the codes, Authors 1 and 5 first coded subsamples of comments independently, and then met to design a codebook, coordinating to add any additional codes as necessary while coding the entire dataset. We categorize codes into key thematic areas—academic, nonacademic, and relationship-building—further refining these with subcategories and attention to whether the advisor take a directive or exploratory approach.
We note that, while these codes provide insight into advisor styles and the content of meetings, they are necessarily limited to what advisors chose to document, and may not fully capture the scope or nuance of each interaction. Details on our coding process are provided in Appendix A.5. Full code definitions are provided in Appendix A.4.1. The open coding process in the first round of analysis familiarized us with key themes in advisor-student interactions. This analysis also serves as a check on data quality, showing whether advisors’ written comments aligned with the interventions they reported taking.
5.1.2 Coding for Expert Targeting
We perform a systematic qualitative analysis of the advisors’ freeform comments to investigate the potential mechanisms behind significant result for the ExpertTest* (§4.2). We focus on comments of meetings where advisors reported taking at least one of the five statistically significant interventions. These codes of advisor-provided reasons enable us to assess whether the observed dependence between action and outcome plausibly reflect non-algorithmic information or these other factors.
Author 1 used a spreadsheet to identify whether advisors explicitly mentioned taking the intervention and provided a reason for doing so (Table 14). For comments that met these requirements, the reasons advisors gave for comments associated with different interventions were identified separately for each intervention, and then merged into a standardized codebook across interventions. We evaluate whether these codes indicate (A) the presence of non-algorithmic, outcome-relevant information guiding the choice of intervention, (B) evidence of targeting at-risk students, or (C) the presence of nuisance confounders (e.g., student requests), in accordance with the questions arising from our quantitative analysis. Full code definitions and examples are provided in Appendix A.4.1.
Recommending a schedule (Intervention 17) was taken as a default action for most students, which, as discussed in §3.1, trivializes the targeted action component of the expertise definition. Thus we additionally code for customization, identifying whether advisors tailored the content of their recommendations based on student-specific context.
@@ 5.1.2 Coding for Expert Targeting
5.2 Qualitative Findings on Non-Algorithmic Targeting
We use the qualitative coding process described above to answer three important questions about the mechanisms for a significant result in the ExpertTest*, as described in §4.2.3: (A) whether non-algorithmic targeting is present, (B) whether advisors support at-risk students in case of the intevention predicting negative outcomes, and (C) whether there is evidence of nuisance confounding. We include selected excerpts from advisors’ comments for illustration; comments are paraphrased as needed for clarity and to preserve anonymity. [^25] To protect student privacy, all gendered pronouns in comments have been changed to they/them.
(A) Advisors use diverse, detailed non-algorithmic context to target interventions.
Across significant interventions, we see a pattern: advisors incorporate non-algorithmic context into their decisions about when and how to intervene. For example, in Intervention 2 (advising a correction to a student’s schedule), advisors incorporate complex information, such as students’ course preferences and history, in their decision to suggest a schedule correction. In one case, an advisor recommends a student postpone a difficult course until the summer based on the student’s self-reflection and planning:
We note that advisors do not always rely on the student’s preference and, in some cases, use their discretion to provide recommendations against the student’s own preference:
“Student needs to add more credits to fall. Offered [MATH XXX]. Student didn’t want to take [MATH XXX] because they said they needed a ‘break’ from math. However, this will help the student stay on track, so I strongly encouraged them to take the class in the fall. Student says they will do so. ”
Unlike the first student, this second student needs to enroll in more courses and does not offer a clear reason for postponing the math course. In this case, the advisor encourages the student to take the class. Advisors, then, not only use non-algorithmic information to decide when to intervene, but also how to intervene, adjusting their particular recommendation—in this case, whether to advise postponing a course—based on context.
These customizations are especially crucial for schedule recommendations (Intervention 17), for which advisors consider detailed factors like student health conditions, motivation to stay in college, and plans to transfer in order to craft schedules for their students:

### v3

@@ 5.1.1 Application to MAAPS Data
The ExpertTest* is framed in terms of individual actions at a single timestep $t$ affecting outcomes at $t+1$, so to the multistage, multi-intervention nature of collegiate advising complicates variable definition. We run different variants of the ExpertTest* to accommodate these complexities. Below we describe how we define $\mathbf{X}_{t}$, $A_{t}$, $Y_{t+1}$, and $n$ for these tests. All test parameters and variables, as well as details of how we chose test parameters $L$ and $K$, are summarized in Appendix A.5.3.
@@
For all the tests, we take $\textbf{X}_{t}$, the algorithmic information, as students’ attributes at baseline excluding their race and ethnicity, since Georgia State University excludes this information from all its early warning systems [77]. Following our data cleaning (§3.1.3), the population consists of 424 students having 3,507 meetings over four years (Appendix Figure 6).
The test requires matching students with approximately identical algorithmically-available features. We pair students greedily using a Euclidean distance function applied to fixed-range scaled features (Appendix Figure 7).We select the number of pairs for each test, $L$, to yield an acceptable tradeoff between test power, which increases in $L$, and test error, which increases in match distance (Appendix A.5.3).
@@
As discussed in §5.1, we also run each test on $A_{t}^{i-}=|1-A_{t}^{i+}|$, which detects cases in which taking the action predicts negative outcomes.
@@
Of the three outcomes we examine, persistence, graduation and GPA (at various timescales), all but GPA are naturally binary; we binarize the GPA outcome by testing for GPA above the mean of 3.5, testing more thresholds for robustness. [^18] We note that the MSE here differs from its normal interpretation—whereas MSE is normally a comparison of predictions of an outcome to that outcome, here we look at whether actions predict binary outcomes. Future work might explore alternate test statistics or conditional independence tests from Alur et al. [6]’s.
@@
    Semesterly tests: Among enrolled students in each semester (Appendix Figure 6), we look at the result of the application of each intervention $i$ in a meeting that semester on term GPA or persistence to the next term (for spring terms, persistence is defined as enrollment/graduated status in either the summer or fall). In these tests, students’ semester of enrollment is added as a feature to their baseline attributes $\mathbf{X}_{t}$, and cross-semester matches of students are prohibited.
@@
Conditional Independence Test.
We test the null
$H_{0}:A_{t}^{i}\perp Y_{t+1}|\ \mathbf{X}_{t}$
for each action $i$ (with both $A_{t}^{i+}$ and $A_{t}^{i-}$), timescale $t$, and outcome $Y$ described above.
@@
5.2 Results
Following correction for multiple testing, four interventions demonstrate significant conditional dependence with term GPA (Table 1) and one intervention demonstrates significant dependence with persistence (Table 2). Significance levels for all 20 interventions are given in Appendix A.5.5, alongside robustness checks with different choices of threshold for GPA, and with exclusion of Advisor 1, who has inconsistent quality of intervention logs.
These five significant interventions comprise 5,615 of the 8,385 total interventions taken by advisors following our data cleaning process. Accordingly, we estimate that 67%, or two-thirds, of interventions taken by MAAPS advisors are potentially expertly targeted. We provide this numerical estimate of the pervasiveness of significant interventions, with the caveat that a significant result is necessary but insufficient for expertly targeted action (which we address by providing qualitative evidence of expertise in §6).
Two of these significance results have a positive direction, indicated by a $(+)$ in Tables 1 and 2, with the $p$-value interpreted as the probability that actions taken by advisors that cannot be explained by administrative data predict a positive outcome for the student (high GPA or next-term enrollment) just by chance. The rest have a negative direction, notated by ()(-), with $p$-values interpreted as the probability that actions predict negative outcomes just by chance. Negative association can be explained by selection bias for students with non-algorithmic risk factors, who may be less likely to succeed on average (see §5.4).
We note that the tests of first-year interventions on a four-year timescale do not meet the significance threshold (see results in Appendix A.5.5). We suspect this is due to the long time range between cause and effect: Advisors’ actions in the first year are unlikely to contain detectable signal about outcomes three years later.
GPA $\geq 3.5$: top five interventions
@@ 5.2 Results
Table 1: Top five most significant interventions for the ExpertTest* for term GPA $\geq 3.5$ (the mean student GPA).
Persistence: top five interventions
@@
Table 2: Top five most significant interventions for the ExpertTest* for persistence to the next term.
5.3 Discussion: Test Power
Interestingly, there is a higher number of significant interventions for high term GPA than for persistence. We hypothesize two primary reasons for this trend. First, the effects of actions advisors take to promote persistence may surface over longer time ranges, whereas semester GPA may be more easily impacted by advisors’ recommendations in a particular term (like recommendations to take particular courses). Second, the power of the ExpertTest* is decreased for outcomes with high or low base rates. Since the test generates synthetic datasets that “swap” intervention values for similar students and compares them to outcomes with the mean squared error, for rare outcomes, there are fewer swaps that can affect the mean squared error of a synthetic dataset, as most individuals’ outcome values are already equal (Appendix A.5.4).
@@ 5.3 Discussion: Test Power
Just like with outcomes, the ExpertTest* may not identify expertly targeted interventions that are very rare. Accordingly, Appendix Table 14 shows that rare interventions in the MAAPS dataset have few swaps between algorithmically-indistinguishable pairs that could change mean squared error, reducing power for tests of these interventions.
5.4 Discussion: Alternative Explanations and Link to Qualitative Analysis
@@
As our Theoretical Framework shows (§4.3), conditional dependence of actions and outcomes is only a necessary condition for expertise and the five significant interventions need not be truly expertly targeted. Conditional dependence of outcomes and actions can have alternative explanations than expertly targeted action.
This fundamental limitation of quantitative analysis underscores the value of our mixed-methods approach. While statistical tests alone cannot definitively establish expertise, qualitative analysis of the advisor meeting notes can complement our understanding of the decision-making process.
In §6, we use qualitative data to provide evidence for the second component of expertly targeted action, called targeted action (Definition 4.1.2), which requires that advisors use non-algorithmic information to decide what action to take. We examine whether comments for significant interventions show evidence of targeting based on non-algorithmic context. For significant interventions that are negatively associated with the outcome, we additionally look for evidence they are targeted based on non-algorithmic risk factors, which place the student at risk for worse outcomes.
Finally, to the extent possible, we look for evidence for and against three alternative explanations of significant results:
Alternative Explanation 1. Nuisance Confounding.
Nuisance confounding between outcomes and actions could contribute to or entirely explain a significant result. For instance, if $A_{t}$ is an advisor making a schedule recommendation, and $Y_{t+1}$ is next-term enrollment, their relationship may be confounded by students requesting scheduling advice—since students who request such advice are disproportionately likely to re-enroll. [^20] Formally, this is a violation of Assumption 4.2, the absence of hidden confounding.
Alternative Explanation 2. Absent Targeting.
Even without nuisance confounding, a significant result still does not imply the presence of targeting. The action may be untargeted, with dependence explained by the action’s effect on the outcome. [^21] Formally, this corresponds to the failure of the targeted action (Definition 4.1.1) component of the expertly targeted action definition.
Alternative Explanation 3. Ineffective Action or Lack of Heterogeneous Effects.
To qualify as expertly targeted, an intervention must actually help students to whom it is targeted. Yet a significant intervention may be targeted based on non-algorithmic information but either harm or fail to affect the outcomes of the targeted student. Alternatively, the action may impact students positively, but not in virtue of the targeting—for example, by helping all students equally. [^22] Formally, these possibilities correspond to the failure of the effective action (Definition 4.1.1) and heterogeneous effects (Definition 4.1.3) components of the expertly targeted action definition, respectively.
@@ 6 Qualitative Analysis: Evidence for Non-Algorithmic Targeted Action
We conduct a two-stage qualitative analysis to better understand advisor discretion and intervention practices, complementing our quantitative findings. The first stage involves open coding to identify themes across all advisor comments, to familiarize ourselves with the data and its limitations. The second stage focuses on a targeted analysis of comments linked to statistically significant interventions from the ExpertTest* to investigate potential evidence of the targeted action component of expertly targeted action (Definition˜4.1). In this section, we describe the methods for our qualitative analysis in §6.1 and how they complement our quantitative findings about advisor expertise in §6.2. In §6.3, we discuss results from the exploratory qualitative analysis in connection to a heterogeneous treatment effect analysis in advisor assignment.
@@ 6.1.1 Exploratory Coding
In the first stage of our qualitative analysis, we conduct open coding to identify themes across all advisor comments and develop familiarity with the data. Following removal of students due to our data cleaning (§3.1.3), we code a total of 1,661 comments left by advisors on meetings with 422 students in the treatment arm. Using the Atlas.ti software, Authors 1 and 5 apply codes for discussion of academic matters (such as course selection and academic performance), nonacademic matters (such as financial aid or family circumstances), and relationship-building matters (such as advisors’ observations on student affect or expressions of empathy for student circumstances).
We note that, while these codes provide insight into advisor styles and the content of meetings, they are necessarily limited to what advisors chose to document, and may not fully capture the scope or nuance of each interaction. Details on our coding process and codebook definitions are provided in Appendix B.1.
6.1.2 Coding for Expertise
Our second round is a targeted analysis of comments to investigate potential evidence of expert targeting. We focus this stage on comments of meetings where advisors reported taking at least one of the five statistically significant interventions (§5.2). Using a spreadsheet, Author 1 first determines whether advisors explicitly mention the relevant intervention and report a clear reason for taking the intervention, which qualifies the comment for inclusion in the analysis. For recommending a schedule (Intervention 17), we additionally code for customization, identifying whether advisors tailored the content of their recommendations based on student-specific context. These criteria are rigidly enforced to mitigate the risk of inferring non-algorithmic targeting due to confirmation bias. For comments meeting the criteria, Author 1 then codes the stated reasons motivating each intervention inductively and merges them into standardized codes. These codes of advisor-provided reasons are assessed for evidence of plausibly non-algorithmic factors informing intervention targeting. Full code definitions and examples are provided in Appendix B.3.1.
@@ 6.1.2 Coding for Expertise
6.2 Qualitative Analysis of Expertly Targeted Action
6.2.1 Evidence of Targeted Action for All Significant Interventions
Int 2 (-) Int 4 (-) Int 16 (-) Int 17 (++) Int 20 (++)
Correct Schedule Degree Progression Schedule Meeting Recommend Schedule Degree Map
1 Course Credits/Placement 3 4 - - 1
2 Course Problem 11 6 6 - 1
3 Degree Progress - 8 - - 2
4 Load Management 18 - - - -
5 Major Choice - 8 18 - -
6 Personal Circumstances - 2 5 - -
7 Low Engagement - - 109 - -
8 Scholarship Process 2 - 1 - -
9 Student-Initiated 3 - 8 $243^{(*)}$ 1
10 Transfer/Dropout - 3 3 - -
11 Unregistered - - 12 - -
Table 3: Reasons found in advisor comments associated with each of the five significant interventions of the ExpertTest*. We indicate how frequently each reason was mentioned by advisors for applying a particular intervention. Most interventions show diverse reasons behind their application. Additionally, interventions associated with negative outcomes, noted with a ()(-), all show evidence of targeting based on risk factors—such as Course Problem, Degree Progress, Load Management, Low Engagement, Transfer/Dropout, and Unregistered (defined in Appendix B.3). Note the exception of Intervention 17: whenever an explicit reason for Recommending a Schedule was mentioned, it was recorded as Student Initiated; however, advisors often incorporated non-algorithmic context into the contents of the schedule (found in 81 comments; SI B.3.3).
We use the qualitative coding procedure described above to determine whether advisors target significant interventions to the student based on non-algorithmic information. We include selected excerpts from advisors’ comments for illustration. [^24] Comments are paraphrased as needed for clarity and to preserve anonymity. To protect student privacy, all gendered pronouns in comments have been changed to they/them.
Across significant interventions, we see a pattern: advisors incorporate non-algorithmic context into their decisions about when and how to intervene. [^25] As discussed further in §6.2.2, Intervention 20 (being advised through in-person development of the degree map/planner) is an exception. We note that the absence of comments demonstrating targeting does not necessarily imply a lack of expert targeting. For example, in providing schedule recommendations (Intervention 17), which is positively associated with persistence ($p=0.039$), advisors consider diverse factors like student health conditions, motivation to stay in college, and plans to transfer to customize their particular schedule recommendations: